# Reinforcement Unlearning via Group Relative Policy Optimization

### v1

@@ Abstract
During pretraining, LLMs inadvertently memorize sensitive or copyrighted data, posing significant compliance challenges under legal frameworks like the GDPR and the EU AI Act. Fulfilling these mandates demands techniques that can remove information from a deployed model without retraining from scratch. Today’s unlearning approaches attempt to address this need, but often leak the very data they aim to erase, sacrifice fluency and robustness, or depend on costly external reward models. We introduce PURGE (Policy Unlearning through Relative Group Erasure), a novel method grounded in the Group Relative Policy Optimization framework that formulates unlearning as a verifiable problem. PURGE uses an intrinsic reward signal that penalizes any mention of forbidden concepts, allowing safe and consistent unlearning. Our approach achieves up to $\times$46 lower token usage per target than state-of-the-art methods, while improving fluency by +5.48% and adversarial robustness by +12.02% over the base model. Extensive evaluation on the Real World Knowledge Unlearning (RWKU) benchmark shows that PURGE reaches 11% unlearning effectiveness while preserving 98% of original utility. PURGE shows that framing LLM unlearning as a verifiable task, enables more reliable, efficient, and scalable forgetting, suggesting a promising new direction for unlearning research that combines theoretical guarantees, improved safety, and practical deployment efficiency.
@@ 1 Introduction
Large Language Models (LLMs) have demonstrated an exceptional capacity to absorb and retain vast amounts of information from large-scale internet datasets. Yet, this capability also brings significant challenges, including the unintentional exposure of sensitive personal data, potential copyright violations, and the risk of misuse in harmful applications. These concerns have gained increasing attention as regulatory frameworks evolve. In the EU, the GDPR (European Union, 2016) establishes the “right to be forgotten”, allowing individuals to request the deletion of their personal data, while the more recent EU AI Act (European Union, 2023) extends these obligations to AI systems, requiring mechanisms to delete specific data on demand. As a result, a growing area of research in AI focuses on techniques to selectively “unlearn” or remove specific information from LLMs without the need for full model retraining. The key challenge is to ensure that this process preserves the model’s overall utility for general downstream language tasks while satisfying regulatory requirements. LLM unlearning has a broad range of applications: it enables privacy compliance by removing memorized sensitive data, supports copyright enforcement by preventing models from reproducing proprietary content, and enhances safety alignment by reducing harmful behaviors (Li et al., 2024) or mitigating biases present in the training data.
@@
Existing LLM unlearning approaches generally fall into three categories. In-context methods use specialized prompts or context manipulations, therefore risking data leakage and consuming limited context-window space. Gradient-ascent fine-tuning can erase memorized data but, if applied too aggressively, often causes model collapse, degrading fluency and utility. Preference-optimization methods treat unlearning as a reward-maximization problem but depend on external reward models, adding overhead computation and complexity. These limitations point to the need for a unified, self-contained approach that (1) prevents residual leakage, (2) preserves model performance, and (3) avoids reliance on external reward models. The shortcomings of these approaches motivate exploring fundamentally different training paradigms. One promising direction is Reinforcement Learning (RL), which has already proven effective in refining LLM behaviors through human feedback. While Reinforcement Learning from Human Feedback (RLHF) has boosted response quality, it remains constrained by noisy annotations and vulnerabilities to reward hacking. In response, DeepSeek (Shao et al., 2024) introduced Reinforcement Learning with Verifiable Rewards (RLVR) (Lambert et al., 2024), a framework that separates tasks into verifiable (e.g., math problem solving, code generation) and non-verifiable (e.g., creative writing, summarization) domains. In verifiable settings, LLMs can repeatedly refine their outputs against objective, measurable criteria, unlocking far greater gains than traditional RLHF.
@@
In this paper, we extend these ideas to the challenge of unlearning in LLMs and introduce PURGE [^1] Code: https://github.com/strzar/purge (Policy Unlearning through Relative Group Erasure), a simple and principled method built on the Group Relative Policy Optimization (GRPO) framework. We treat unlearning as a verifiable task, where the successful removal of specific data can be directly measured. PURGE builds a reward function based on clear, measurable indicators of data removal, allowing the model to optimize forgetting the same way Large Reasoning Models (LRMs) optimize reasoning. Our results show that this verifiable RL approach is more reliable than existing unlearning techniques and provides a scalable, accountable solution for safely removing information from LLMs.
Specifically, our contributions are: (1) Principled Unlearning Framework – We propose PURGE, an RL-based approach that treats unlearning as a verifiable task. Unlike prior methods that try to remove specific data, PURGE leverages GRPO to guide LLMs in forgetting specific knowledge while maintaining general utility. (2) Theoretical Guarantees – We provide formal results on the suppression of targeted knowledge, proving geometric decay of forbidden-token probabilities and high-probability bounds on utility retention via KL divergence. (3) Scalable and Efficient Unlearning – PURGE achieves competitive unlearning performance using significantly fewer tokens – up to $\times 46$ fewer per forget target compared to SoTA approaches – while requiring no external reward models, making it a more scalable and cost-effective approach for real-world deployments. (4) Comprehensive Empirical Evaluation – Extensive experiments on the RWKU (Jin et al., 2024) benchmark – spanning knowledge memorization, knowledge manipulation, adversarial robustness and real-world utility tasks – validate that PURGE achieves more natural and coherent outputs, improving fluency by +5.48% compared to the base model, and showing +12.02% greater resistance to adversarial attacks, ensuring safer and more reliable unlearning.
@@ 2.1 What is Machine Unlearning?
For notational clarity we treat the input and output spaces as distinct, even though in our setting they coincide – i.e., $\mathcal{X}=\mathcal{Y}=\mathcal{V}$, where $\mathcal{V}$ is the set of all finite token sequences. We simply use $\mathcal{X}$ to emphasize “input” and $\mathcal{Y}$ to emphasize “output.” We consider a parametric model family
@@
where $\ell\colon\mathcal{Y}\times\mathcal{Y}\to\mathbb{R}_{\geq 0}$ denotes the per-example loss measuring discrepancy between prediction and true label. This ensures that, in expectation, the unlearned model generalizes to unseen data just as well as the original.
@@ 2.2 LLM Unlearning via Empirical Proxies
$\widehat{R}_{R}(\theta)=\frac{\underset{(x,y)\in\mathscr{D}_{R}}{\sum}\ell(f_{\theta}(x),y)}{|\mathscr{D}_{R}|},\;|\widehat{R}_{R}(\theta^{\prime})-\widehat{R}_{R}(\theta^{*})|\leq\epsilon_{R},$ (7)
@@
$\widehat{R}_{T}(\theta)=\frac{\underset{(x,y)\in\mathscr{D}_{T}}{\sum}\ell(f_{\theta}(x),y)}{|\mathscr{D}_{T}|},\;|\widehat{R}_{T}(\theta^{\prime})-\widehat{R}_{T}(\theta^{*})|\leq\epsilon_{G},$ (8)
with a tolerance $\epsilon_{G}\geq 0$. These two high-confidence, distribution-free checks serve as operational proxies for the retention and generalization desiderata without assuming knowledge of $\mathscr{D}$ or the underlying data distribution. All LLM-specific evaluations – e.g., token likelihood suppression, alignment scores, privacy leakage tests, downstream task metrics – are carried out within this empirical framework. We provide the proof of the equivalence of empirical proxies to the original conditions in the Appendix.
@@ 3.1 Machine Unlearning
Earlier work on machine unlearning (MU) focused on classical machine learning, using influence functions and exact deletion methods to manage data effects (Cao and Yang, 2015; Hoofnagle et al., 2019; Bourtoule et al., 2021; Nguyen et al., 2022). This capability has driven its adoption in various domains. In image classification, approaches like certified data removal and adaptive retraining guarantee that deleted samples have negligible impact on predictions (Ginart et al., 2019; Golatkar et al., 2020; Neel et al., 2021; Ullah et al., 2021; Sekhari et al., 2021). More recently, MU has been integrated into text-to-image frameworks to limit the spread of copyrighted or sensitive content (Fan et al., 2024). Federated learning has also benefited, enabling individual clients to withdraw their contributions while preserving model utility and privacy (Liu et al., 2020; Che et al., 2023; Halimi et al., 2022). Finally, extensions to graph neural networks address unlearning at the node and edge level, ensuring minimal residual influence on downstream tasks (Chen et al., 2022; Chien et al., 2022; Wu et al., 2023).
@@ 3.2 Unlearning in LLMs
The unique challenges of large language models (LLMs) like their vast size and black-box nature have made unlearning a critical research area (Blanco-Justicia et al., 2025; Liu et al., 2024; 2025; Geng et al., 2025). Current unlearning methods for LLMs fall into two main categories: direct parameter updates and preference optimization frameworks with complementary work proposing robust and parameter-efficient unlearning through unified or low-overhead fine-tuning schemes that mitigate collateral forgetting (Cha et al., ; Ding et al., ).
Direct Fine-Tuning Strategies. These methods directly modify the model’s parameters. Relabeling fine-tuning replaces unwanted outputs in a “forget set” with neutral or refusal responses and then uses standard gradient descent to overwrite the original knowledge (Eldan and Russinovich, 2023). Gradient Ascent (GA), on the other hand, maximizes the next-token loss on the forget set, actively pushing the model away from the knowledge to be forgotten (Jang et al., 2023). Because pure GA can severely degrade a model’s fluency, many approaches add retain-set regularizers – either a concurrent gradient-descent loss on retained data (Liu et al., 2022) or a KL-divergence penalty that keeps the updated model close to its original distribution (Yao et al., 2024). Several studies analyze the objectives and gradient dynamics underlying loss-based unlearning (Wang et al., ; Wang et al., ; Yuan et al., ), highlighting failure modes of existing approaches and proposing more principled formulations.
Preference-Optimization Methods. These methods reframe unlearning as a preference task. Quark uses reinforcement learning (RL) to unlearn, penalizing undesirable outputs with a reward model while using a KL term to preserve the model’s style (Lu et al., 2022). DeMem extends this with a negative-similarity reward that encourages the model to paraphrase while retaining original meaning (Kassem et al., 2023). Direct Preference Optimization (DPO) adapts unlearning into a binary preference task where refusal answers are preferred over unwanted outputs (Rafailov et al., 2023). Negative Preference Optimization (NPO) simplifies this by only using negative examples, and it has been shown to converge more smoothly than GA (Zhang et al., 2024). A common baseline is Rejection Tuning (RT), which trains the model to refuse queries related to the forget set by fine-tuning on data where the target answer is replaced with a fixed refusal template like “I don’t know” (Jin et al., 2024; Maini et al., 2024). While RT is effective at suppressing outputs, it can create shortcuts and leave latent traces that may re-emerge under certain conditions.
@@ 4.2 Utilizing GRPO
We adapt the GRPO (Shao et al., 2024) algorithm for PURGE (see Algorithm 1) to fine-tune the policy so that it reliably unlearns a specified vocabulary while maintaining fluency and task performance. GRPO is a Direct Policy Optimization (DPO) variant of PPO that (i) compares multiple candidate answers to the same prompt, (ii) computes group-relative advantages, and (iii) adds a KL term directly to the loss instead of to the reward. These modifications remove the need for an external reward network and align naturally with reward models that are themselves trained on pairwise comparisons.
@@ 4.3 Theoretical Analysis
Gist of Theorem 2: Because GRPO includes an explicit KL‐penalty tying the updated policy back to the original, any drop in downstream performance on the retained tasks is at most on the order of the square root of that KL divergence. In practice, this means if you choose a moderate KL weight, you can almost entirely preserve your model’s utility while safely removing forbidden content.
@@ 5.1 Experimental Setup
We compare PURGE against GA (Jang et al., 2023), DPO (Rafailov et al., 2023), NPO (Zhang et al., 2024), and RT (Jin et al., 2024; Maini et al., 2024), across four benchmark dataset splits by relying on the RWKU benchmarking framework. Each dataset split employs a distinct evaluation metric to probe a different aspect of our fine‐tuned model’s unlearning capability. In addition, we report two baselines commonly found in the literature: (1) the performance of our base model without any unlearning procedure, and (2) In‐Context Unlearning (ICU) (Thaker et al., 2024; Pawelczyk et al., 2024). We conduct our main experiments using Phi-3-Mini-4K-Instruct; a 3.8 billion parameter model and we utilise the default hyperparameters for the SoTA unlearning methods. For PURGE, we first generate the synthetic probes training dataset and the $\mathscr{D}^{\prime}_{F}$ for each unlearning target. Following RWKU’s evaluation process, we run each unlearning request independently from each other unlearning target and we average all the runs. We perform our experiment on one AMD EPYC 7002/3 64-Core CPU and one Nvidia TESLA A100 GPU, with a total of $\sim$350hh execution time.
@@ 5.2 Results
Table 1 reports the performance of five unlearning methods and two baselines on the RWKU Famous People dataset, evaluated on forgetting quality, neighboring knowledge recall, membership‑privacy resistance, and overall utility. PURGE achieves substantial reductions in recall on the Forget Set, with a $\Delta$(us, base) of 5.53% in the Fill‑in‑the‑Blank (FB) split, 10.07% in the Question‑Answering (QA) split, and 12.02% in the Adversarial Attack (AA) split. Similarly, it preserves neighboring knowledge comparable to SoTA. In terms of privacy attacks, PURGE attains an FM score of 40.26 and an RM score of 38.64 (tied for best), demonstrating effective unlearning while maintaining robustness against Membership Inference Attacks (MIA). Finally, PURGE sustains downstream task utility, trailing slightly behind SoTA. This small gap is expected given the inherent trade‑off between unlearning strength, utility, and the reported standard deviations. Although NPO outperforms PURGE in overall utility – owing to its more comprehensive forget corpus, which in real world scenarios is difficult to aquire, as mentioned in section 4.1 – PURGE remains competitive. Moreover, PURGE achieves the highest fluency score (FLU = 140.90), validating our hypothesis that guiding the model via GRPO based unlearning yields more natural and coherent outputs, a challenge for many existing approaches.
Table 1: Unlearning with Phi-3-Mini-4K-Instruct on the RWKU Famous People Dataset. PURGE outperforms the Base (intact) model on 8/12 aspects, ranking second-best on 5/12 and first on 2/12. ✓ indicates the corresponding theorem is empirically satisfied. * denotes no utility loss (performance $\geq$ original). Bold marks the best overall average; underline, second-best. FM, RM, and FLU are reported as sums per RWKU benchmark.

### v2

@@ Abstract
During pretraining, LLMs inadvertently memorize sensitive or copyrighted data, posing significant compliance challenges under legal frameworks like the GDPR and the EU AI Act. Fulfilling these mandates demands techniques that can remove information from a deployed model without retraining from scratch. Today’s unlearning approaches attempt to address this need, but often leak the very data they aim to erase, sacrifice fluency and robustness, or depend on costly external reward models. We introduce PURGE (Policy Unlearning through Relative Group Erasure), a novel method grounded in the Group Relative Policy Optimization framework that formulates unlearning as a verifiable problem. PURGE uses an intrinsic reward signal that penalizes any mention of forbidden concepts, allowing safe and consistent unlearning. Our approach achieves up to $\times$46 lower token usage per target than state-of-the-art methods, while improving fluency by +5.48% and adversarial robustness by +12.02% over the base model. Extensive evaluation on the Real World Knowledge Unlearning (RWKU) benchmark shows that PURGE reaches 11% unlearning effectiveness while preserving 98% of original utility. PURGE shows that framing LLM unlearning as a verifiable task enables more reliable, efficient, and scalable forgetting, suggesting a promising new direction for unlearning research that combines theoretical guarantees, improved safety, and practical deployment efficiency.
@@ 1 Introduction
Large Language Models (LLMs) have demonstrated an exceptional capacity to absorb and retain vast amounts of information from large-scale internet datasets. Yet this capability also poses significant challenges, including unintentional exposure of sensitive personal data, potential copyright violations, and the risk of misuse in harmful applications. These concerns have gained increasing attention as regulatory frameworks evolve. In the EU, the GDPR (European Union, 2016) establishes the “right to be forgotten”, allowing individuals to request the deletion of their personal data. In contrast, the more recent EU AI Act (European Union, 2023) extends these obligations to AI systems, requiring mechanisms to delete specific data on demand. As a result, a growing area of AI research focuses on techniques to selectively “unlearn” or remove specific information from LLMs without the need for full model retraining. The key challenge is to ensure that this process preserves the model’s overall utility for general downstream language tasks while satisfying regulatory requirements. LLM unlearning has a broad range of applications. For example, it enables privacy compliance by removing memorized sensitive data, supports copyright enforcement by preventing models from reproducing proprietary content, and enhances safety alignment by reducing harmful behaviors (Li et al., 2024) or mitigating biases present in the training data.
@@
Existing LLM unlearning approaches generally fall into three categories. In-context methods use specialized prompts or context manipulations, therefore risking data leakage and consuming limited context-window space. Gradient-ascent fine-tuning can erase memorized data, but if applied too aggressively, it often causes model collapse, degrading fluency and utility. Preference-optimization methods treat unlearning as a reward-maximization problem but rely on external reward models, which add overhead in computation and complexity. These limitations point to the need for a unified, self-contained approach that (1) prevents residual leakage, (2) preserves model performance, and (3) avoids reliance on external reward models. The shortcomings of these approaches motivate the exploration of fundamentally different training paradigms. One promising direction is Reinforcement Learning (RL), which has already proven effective at refining LLM behavior with human feedback. While Reinforcement Learning from Human Feedback (RLHF) has improved response quality, it remains constrained by noisy annotations and is vulnerable to reward hacking. In response, DeepSeek (Shao et al., 2024) introduced Reinforcement Learning with Verifiable Rewards (RLVR) (Lambert et al., 2024). This framework separates tasks into verifiable (e.g., math problem solving, code generation) and non-verifiable (e.g., creative writing, summarization) domains. In verifiable settings, LLMs can repeatedly refine their outputs against objective, measurable criteria, unlocking far greater gains than traditional RLHF.
@@
In this paper, we extend these ideas to the challenge of unlearning in LLMs and introduce PURGE (Policy Unlearning through Relative Group Erasure), a simple and principled method built on the Group Relative Policy Optimization (GRPO) framework. We treat unlearning as a verifiable task, with the successful removal of specific data directly measurable. PURGE builds a reward function based on clear, measurable indicators of data removal, allowing the model to optimize forgetting the same way Large Reasoning Models (LRMs) optimize reasoning. Our results show that this verifiable RL approach is more reliable than existing unlearning techniques and provides a scalable, accountable solution for safely removing information from LLMs.
Specifically, our contributions are:
(1) Principled Unlearning Framework. We propose PURGE, an RL-based approach that treats unlearning as a verifiable task. Unlike prior methods that aim to remove specific data, PURGE leverages GRPO to guide LLMs to forget specific knowledge while maintaining general utility.
(2) Theoretical Guarantees. We provide formal results on the suppression of targeted knowledge, proving geometric decay of forbidden-token probabilities and high-probability bounds on utility retention via KL divergence.
(3) Scalable and Efficient Unlearning. PURGE achieves competitive unlearning performance with significantly fewer tokens – up to $\times 46$ fewer per forget target than SoTA approaches – while requiring no external reward models, making it a more scalable and cost-effective approach for real-world deployments.
(4) Comprehensive Empirical Evaluation. Extensive experiments on the RWKU (Jin et al., 2024) benchmark – spanning knowledge memorization, knowledge manipulation, adversarial robustness and real-world utility tasks – validate that PURGE achieves more natural and coherent outputs, improving fluency by +5.48% compared to the base model, and showing +12.02% greater resistance to adversarial attacks, ensuring safer and more reliable unlearning.
@@ 2.1 What is Machine Unlearning?
For notational clarity, we treat the input and output spaces as distinct, even though in our setting they coincide – i.e., $\mathcal{X}=\mathcal{Y}=\mathcal{V}$, where $\mathcal{V}$ is the set of all finite token sequences. We simply use $\mathcal{X}$ to emphasize “input” and $\mathcal{Y}$ to emphasize “output.” We consider a parametric model family
@@
where $\ell\colon\mathcal{Y}\times\mathcal{Y}\to\mathbb{R}_{\geq 0}$ denotes the per-example loss measuring discrepancy between prediction and true label. This ensures that, in expectation, the unlearned model generalizes to unseen data as well as the original model.
@@ 2.2 LLM Unlearning via Empirical Proxies
$\widehat{R}_{R}(\theta)=\frac{\underset{(x,y)\in\mathscr{D}_{R}}{\sum}\ell(f_{\theta}(x),y)}{|\mathscr{D}_{R}|},\qquad|\widehat{R}_{R}(\theta^{\prime})-\widehat{R}_{R}(\theta^{*})|\leq\epsilon_{R},$ (7)
@@
$\widehat{R}_{T}(\theta)=\frac{\underset{(x,y)\in\mathscr{D}_{T}}{\sum}\ell(f_{\theta}(x),y)}{|\mathscr{D}_{T}|},\qquad|\widehat{R}_{T}(\theta^{\prime})-\widehat{R}_{T}(\theta^{*})|\leq\epsilon_{G},$ (8)
with a tolerance $\epsilon_{G}\geq 0$. These two high-confidence, distribution-free checks serve as operational proxies for the retention and generalization desiderata without assuming knowledge of $\mathscr{D}$ or the underlying data distribution. All LLM-specific evaluations – e.g., token-likelihood suppression, alignment scores, privacy-leakage tests, and downstream task metrics – are conducted within this empirical framework. We provide the proof of the equivalence of empirical proxies to the original conditions in the Appendix.
@@ 3.1 Machine Unlearning
Earlier work on machine unlearning (MU) focused on classical machine learning, using influence functions and exact deletion methods to manage data effects (Cao and Yang, 2015; Hoofnagle et al., 2019; Bourtoule et al., 2021; Nguyen et al., 2025). This capability has driven its adoption in various domains. In image classification, approaches such as certified data removal and adaptive retraining ensure that deleted samples have a negligible impact on predictions (Ginart et al., 2019; Golatkar et al., 2020; Neel et al., 2021; Ullah et al., 2021; Sekhari et al., 2021). More recently, MU has been integrated into text-to-image frameworks to limit the spread of copyrighted or sensitive content (Fan et al., 2024). Federated learning has also benefited, enabling individual clients to withdraw their contributions while preserving model utility and privacy (Liu et al., 2020; Che et al., 2023; Halimi et al., 2022). Finally, extensions to graph neural networks address unlearning at the node and edge level, ensuring minimal residual influence on downstream tasks (Chen et al., 2022; Chien et al., 2022; Wu et al., 2023).
@@ 3.2 Unlearning in LLMs
The unique challenges of LLMs, such as their vast size and black-box nature, have made unlearning a critical research area (Blanco-Justicia et al., 2025; Liu et al., 2024; 2025; Geng et al., 2025). Current unlearning methods for LLMs fall into two main categories: direct parameter updates and preference optimization frameworks, with complementary work proposing robust, parameter-efficient unlearning through unified or low-overhead fine-tuning schemes that mitigate collateral forgetting (Cha et al., 2025; Ding et al., 2025).
Direct Fine-Tuning Strategies. These methods directly modify the model’s parameters. Relabeling fine-tuning replaces unwanted outputs in a “forget set” with neutral or refusal responses and then uses standard gradient descent to overwrite the original knowledge (Eldan and Russinovich, 2023). Gradient Ascent (GA), on the other hand, maximizes the next-token loss on the forget set, actively pushing the model away from the knowledge to be forgotten (Jang et al., 2023). Because pure GA can severely degrade a model’s fluency, many approaches add retain-set regularizers – either a concurrent gradient-descent loss on retained data (Liu et al., 2022) or a KL-divergence penalty that keeps the updated model close to its original distribution (Yao et al., 2024). Several studies analyze the objectives and gradient dynamics underlying loss-based unlearning (Wang et al., 2025b; a; Yuan et al., 2025), highlighting failure modes of existing approaches and proposing more principled formulations.
Preference-Optimization Methods. These methods reframe unlearning as a preference task. Quark uses reinforcement learning (RL) to unlearn, penalizing undesirable outputs with a reward model while using a KL term to preserve the model’s style (Lu et al., 2022). DeMem extends this with a negative-similarity reward that encourages the model to paraphrase while retaining original meaning (Kassem et al., 2023). Direct Preference Optimization (DPO) adapts unlearning to a binary preference task in which refusal answers are preferred over unwanted outputs (Rafailov et al., 2023). Negative Preference Optimization (NPO) simplifies this by only using negative examples, and it has been shown to converge more smoothly than GA (Zhang et al., 2024). A common baseline is Rejection Tuning (RT), which trains the model to refuse queries related to the forget set by fine-tuning on data where the target answer is replaced with a fixed refusal template like “I don’t know” (Jin et al., 2024; Maini et al., 2024). While RT is effective at suppressing outputs, it can create shortcuts and leave latent traces that may re-emerge under certain conditions.
@@ 4.2 Utilizing GRPO
We adapt the GRPO (Shao et al., 2024) algorithm for PURGE (see Algorithm 1) to fine-tune the policy so that it reliably unlearns a specified vocabulary while maintaining fluency and task performance. GRPO is a Direct Policy Optimization (DPO) variant of PPO that (i) compares multiple candidate answers to the same prompt, (ii) computes group-relative advantages, and (iii) adds a KL term directly to the loss instead of to the reward. These modifications eliminate the need for an external reward network and align naturally with reward models trained on pairwise comparisons.
@@ 4.3 Theoretical Analysis
Gist of Theorem 2: Because GRPO includes an explicit KL penalty tying the updated policy back to the original, any drop in downstream performance on the retained tasks is at most on the order of the square root of that KL divergence. In practice, this means that if you choose a moderate KL weight, you can almost entirely preserve your model’s utility while safely removing forbidden content.
@@ 5.1 Experimental Setup
We compare PURGE against GA (Jang et al., 2023), DPO (Rafailov et al., 2023), NPO (Zhang et al., 2024), and RT (Jin et al., 2024; Maini et al., 2024), across four benchmark dataset splits by relying on the RWKU benchmarking framework. Each dataset split employs a distinct evaluation metric to probe a different aspect of our fine‐tuned model’s unlearning capability. In addition, we report two baselines commonly found in the literature: (1) the performance of our base model without any unlearning procedure, and (2) In‐Context Unlearning (ICU) (Thaker et al., 2024; Pawelczyk et al., 2024). We conduct our main experiments using Phi-3-Mini-4K-Instruct, a 3.8-billion-parameter model, and utilise the default hyperparameters for the SoTA unlearning methods. For PURGE, we first generate the synthetic probes training dataset and the $\mathscr{D}^{\prime}_{F}$ for each unlearning target. Following RWKU’s evaluation process, we run each unlearning request independently from each other unlearning target, and we average all the runs. We perform our experiment on one AMD EPYC 7002/3 64-Core CPU and one Nvidia TESLA H200 GPU, with a total execution time of $\sim$350hh.
@@ 5.2 Results
Table 1 reports the performance of five unlearning methods and two baselines on the RWKU Famous People dataset, evaluated on forgetting quality, neighboring knowledge recall, membership‑privacy resistance, and overall utility. PURGE achieves substantial reductions in recall on the Forget Set, with a $\Delta$(us, base) of 5.53% in the Fill‑in‑the‑Blank (FB) split, 10.07% in the Question‑Answering (QA) split, and 12.02% in the Adversarial Attack (AA) split. Similarly, it preserves neighboring knowledge comparable to SoTA. In terms of privacy attacks, PURGE attains an FM score of 40.26 and an RM score of 38.64 (tied for best), demonstrating effective unlearning while maintaining robustness against Membership Inference Attacks (MIA). Finally, PURGE sustains downstream task utility, trailing slightly behind SoTA. This small gap is expected, given the inherent trade‑off between unlearning strength and utility, as reflected in the reported standard deviations. Although NPO outperforms PURGE in overall utility – owing to its more comprehensive forget corpus, which in real-world scenarios is difficult to acquire, as mentioned in section 4.1 – PURGE remains competitive. Moreover, PURGE achieves the highest fluency score (FLU = 140.90), validating our hypothesis that guiding the model via GRPO-based unlearning yields more natural and coherent outputs, a challenge for many existing approaches.
Table 1: Unlearning with Phi-3-Mini-4K-Instruct on the RWKU Famous People Dataset. PURGE outperforms the Base (intact) model on 8/12 aspects, ranking second-best on 5/12 and first on 2/12. ✓ indicates the corresponding theorem is empirically satisfied. * denotes no utility loss (performance $\geq$ original). Bold marks the best overall average; underline, second-best. FM, RM, and FLU are reported as sums per the RWKU benchmark.